The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field.

Answer by checkley75(3666)

L=(2W-5) THEN 2W+2L=230
2W+2(2W-5)=230
2W+4W-10=230
6W=230+10
6W=240
W=240/6
W=40 FT.
L=2*40-5
L=80-5
L=75 FT.
PROOF
2*40+2*75=230
80+150=230
230=230